f you want to be
95
95​% confident of estimating the population proportion to within a sampling error of
plus or minus 0.08 comma
±0.08, what sample size is​ needed?

Answer by Boreal(15235)

Assume p=0.5, which is most conservative
error=z*sqrt*p*(1-p))/n
error^2=z^2*0.25/n
n*err^2=0.25 z^2
n*0.0064=0.25*1.96^2
n=150.0625 or 151